# Uniqueness of the inverse source problem for fractional diffusion-wave equations

### v1

@@ 2 Preliminaries and main theorems
This section extends the discussion to equation (1). Define the domain of $A$ by $\mathcal{D}(A)=H^{2}(\Omega)\cap H_{0}^{1}(\Omega)$. Suppose $f\in\mathcal{D}(A)$ and $\mu\in L^{\infty}(0,\infty)$, then equation (1) admits a weak solution $u\in L^{2}_{\text{loc}}(0,\infty;$ $\mathcal{D}(A))\cap H^{\alpha}_{\text{loc}}(0,\infty;L^{2}(\Omega))$ as established in [19, Theorem 2.2]. Given that the spectrum of $A$ consists of positive eigenvalues of finite multiplicities, these can be denoted as $\{\lambda_{n}\}_{n\in\mathbb{N}}$ in increasing order, without counting multiplicities. Let $P_{n}$ be the orthogonal projection onto the eigenspace of $\lambda_{n}$. The solution $u$ is given by (refer to [22, Lemma 1])
@@
(a) $E_{\alpha,\beta}$ is entire in $\mathbb{C}$;
@@
$E_{\alpha,\beta}(-x)\sim\sum_{k=0}^{\infty}\frac{(-1)^{k+1}}{\Gamma(\beta-k\alpha)}x^{-k}\text{ as }x\rightarrow+\infty;$
@@
Then $||g||_{H^{2\sigma}(\Omega)}\leq C||A^{\sigma}g||_{L^{2}(\Omega)}$ for $g\in\mathcal{D}(A^{\sigma})$. Moreover, we denote $g=O(t^{-\infty})$ if $g=O(t^{-p})$ for any $p\in\mathbb{N}$.
@@
Assume that $\mu$ has an asymptotic expansion over $t^{-1}$:
@@
(a) Assume that $f\in\mathcal{D}(A)$ and
@@
(b) Let $f\in\mathcal{D}(A^{\sigma_{2}})$, where $\sigma_{2}>\max\{1,(2d-1)/4\}$. Assume that
@@
Let $\mu$ satisfy the conditions (3) and (4a)/(4b) in Theorem 2.2. Define
@@
(a) Let $f\neq 0$ and $f\in\mathcal{D}(A^{\sigma_{3}})$, where $\sigma_{3}>\max\{1,3d/4-1\}$. Then for any $x_{0}\in\Omega\setminus\Lambda(f)$,
@@
(b) Let $f\neq 0$ and $f\in\mathcal{D}(A^{\sigma_{4}})$, where $\sigma_{4}>\max\{1,(3d-1)/4\}$. Then for any $x_{0}\in\partial\Omega\setminus\Lambda_{b}(f)$,
@@
Let $\alpha\in(0,2)\setminus\mathbb{Q}$. Suppose $u$ and $\widetilde{u}$ are solutions to the fractional diffusion-wave equation (1) with source terms $F(x,t)=f(x)\mu(t)$ and $\widetilde{F}(x,t)=\widetilde{f}(x)\widetilde{\mu}(t)$ respectively. Moreover, $\mu$ and $\widetilde{\mu}$ satisfy the conditions (3) and (4a)/(4b) in Theorem 2.2.
(a) Assume that $f,\widetilde{f}\in\mathcal{D}(A)$ and
@@
(b) Let $f,\widetilde{f}\in\mathcal{D}(A^{\sigma_{2}})$, where $\sigma_{2}>\max\{1,(2d-1)/4\}$. Assume that
@@ 3.1 The key propositions
Take the last term as a constant dependent on $t$, and the proof is complete. $\square$
@@ 3.2 The asymptotic expansion of $\psi_{n}(t)$
Since $\overline{R}_{\mu,m+1}(s)=O(s^{-m})$ as $s\rightarrow 0$ and $\overline{R}_{\mu,m+1}(s)=O(s^{-m-2})$ as $s\rightarrow\infty$, the quantity $c_{\mu,m}:=\int_{0}^{\infty}s^{m}\overline{R}_{\mu,m+1}(s)\,\text{d}s$ is well defined.
@@
In the framework of Proposition 3.6, for any $N\in\mathbb{N}$, one can find $J=J(N)$, $K=K(N)$, and $M=M(N)$ such that $\psi_{n}(t)$ expands as follows:
@@
Proof. Since $\psi_{n}$ is linear with respect to $\mu$, we first consider the zeroth order term $\mu_{0}$. Note that $E_{\alpha,1}(z)^{\prime}=\alpha^{-1}E_{\alpha,\alpha}(z)$ by definition, so
@@
We emphasize that the bounding constant in the big-O notation is independent of $\lambda_{n}$.
In the remaining proof, we set $\mu_{0}=0$ and introduce two parameters $\beta_{1}$ and $\beta_{2}$ in $(0,1)$. Write $\mu(t)=\sum_{j=1}^{J}\mu_{j}t^{-j}+R_{\mu,J}(t)$, then $|R_{\mu,J}(t)|<Ct^{-J-1}$ for $t>1$. Since
@@
with $R_{E,K}(\lambda_{n}s^{\alpha})=O\big{(}(\lambda_{n}s^{\alpha})^{-l_{K+1}-1}\big{)}$. We can choose $T_{1}$ sufficiently large and independent of $\lambda_{n}$ such that $t-t^{\beta_{1}}>\lambda_{n}^{-1/\alpha}t^{\beta_{2}}$ when $t>T_{1}$. Note that $\mu\in L^{\infty}(0,\infty)$, so
@@
Besides, for each $j$ we have
@@
For $N\in\mathbb{N}$ given above, we choose $K=K(N)$ such that $\alpha\beta_{2}l_{K+1}>N$. Combining equations (13) ~ (16), we obtain
@@
Now we would like to expand the term $(t-s)^{\gamma},\gamma<0$ in the integrals. Write
$(1-s)^{\gamma}=\sum_{m=0}^{M}(-1)^{m}\binom{\gamma}{m}s^{m}+R_{\gamma,M}(s).$
Then
@@
It remains to expand integrals in the form of $\int_{0}^{t^{\beta}}s^{\gamma}F(s)\,\text{d}s$. For the integral $\int_{0}^{t^{\beta_{1}}}s^{m}\mu(s)\,\text{d}s$, we have to use the alternative expansion (10) of $\mu$. By equation (11), $\overline{R}_{\mu,J^{\prime}}=O(t^{-J^{\prime}-1})$ for $J^{\prime}>m$. Hence,
@@
$\int_{0}^{t^{\beta_{1}}}s^{m}\mu(s)\,\text{d}s$ $=\sum_{\begin{subarray}{c}1\leq j\leq J^{\prime}\\ j\neq m+1\end{subarray}}b_{j}^{\times}t^{\beta_{1}(m-j+1)}+\mu_{m+1}\log(1+t^{\beta_{1}})+c_{\mu,m}+O(t^{-N})$ (21)
$=\sum_{\begin{subarray}{c}1\leq j\leq J^{\prime}\\ j\neq m+1\end{subarray}}b_{j}^{\times}t^{\beta_{1}(m-j+1)}+\mu_{m+1}\beta_{1}\log t+c_{\mu,m}+O(t^{-N}).$
Similarly we can expand $\int_{0}^{t^{\alpha\beta_{2}}}s^{m/\alpha}E_{\alpha,\alpha}(-s)\,\text{d}s$. When $m/\alpha\in\mathbb{N}\cup\{0\}$, observe that the coefficient of $s^{-m/\alpha-1}$ in the expansion of $E_{\alpha,\alpha}(-s)$ is 00, so for any $m\in\mathbb{N}\cup\{0\}$, we can use the ordinary expansion of $E_{\alpha,\alpha}$:
@@
$\int_{0}^{t^{\alpha\beta_{2}}}s^{m/\alpha}E_{\alpha,\alpha}(-s)\,\text{d}s=\sum_{k=1}^{K^{\prime}}b_{j,k}t^{\beta_{2}(m-l_{k}\alpha)}+b_{j,m}+O(t^{-N})$ (22)
for some $K^{\prime}=K^{\prime}(N)$.
The remaining term $\int_{\lambda_{n}^{-1/\alpha}t^{\beta_{2}-1}}^{1-t^{\beta_{1}-1}}s^{-l_{k}\alpha-1}(1-s)^{-j}\,\text{d}s$ can be expressed by the incomplete Beta function $B(a,b;x)$ and the hypergeometric function ${}_{2}\mathbf{F}_{1}(a,b,c;x)$ (see section 8.17 and chapter 15 of [16]). By equations (8.17.7) and (15.1.2) of [16] we obtain
@@
Finally we claim that the expansion of $\psi_{n}$ should be independent of $\beta_{1}$ and $\beta_{2}$ due to the uniqueness of asymptotic expansion. We choose $(\beta_{1},\beta_{2})$ and $(\beta_{1},\beta_{2})$ such that $\beta_{1}$, $\beta_{2}$, $\beta_{1}^{\prime}$, $\beta_{2}^{\prime}$ and $\alpha$ are algebraically independent over $\mathbb{Z}$. It follows that, if we pick a term in the first sum of the expansion of $\psi_{n}(t)$ with parameters $(\beta_{1},\beta_{2})$, the term still occurs under parameters $(\beta_{1}^{\prime},\beta_{2}^{\prime})$ if and only if the coefficients of $\beta_{1}$ and $\beta_{2}$ on the exponent of $t$ vanishes. Note that $l_{k}\alpha\notin\mathbb{N}\cup\{0\}$ by definition, so the only remaining term in the first sum is $\sum_{j=1}^{J^{\prime\prime}}\sum_{k=1}^{K^{\prime\prime}}b_{jk,j-1}^{\times}\lambda_{n}^{-l_{k}-1}t^{-l_{k}\alpha-j}$. This term can be brought into the second sum, again due to $l_{k}\alpha\notin\mathbb{N}\cup\{0\}$. $\square$
@@ 3.3 Proof of the key propositions
$\sum_{n=1}^{\infty}|a_{n}\lambda_{n}^{-1}\log(1+\lambda_{n})|<\infty,$
By the decay condition (8), all the coefficients in the expansion above (after collecting the terms) should be zero.
If there exists $j\geq 1$ such that $\mu_{j}\neq 0$, we choose the smallest such $j$ and denote it by $j_{0}$. In this situation we extract terms of the form $t^{\gamma}\log t$ from the expansion (27). When $j=j_{0}$ and $k=1$, the exponent $-l_{k}\alpha-j$ of $t$ is the largest among those with nonzero coefficients. Since $l_{1}\alpha\notin\mathbb{N}\cup\{0\}$, we have $\frac{(-1)^{l_{1}-1+j_{0}}}{\Gamma(-l_{1}\alpha)}\neq 0$, which implies $A_{l_{1}}=0$. Now the summation over $k$ in expansion (27) can start from $k=2$, and the exponent $-l_{k}\alpha-j$ takes the maximum when $j=j_{0}$ and $k=2$. Repeat the process above, and we can prove that $A_{l_{k}}=0$ for all $k\in\mathbb{N}$. Finally the proposition is proved by applying Lemma 3.7 to $\{A_{l_{k}}\}_{k\in\mathbb{N}}$.
If $\mu_{0}\neq 0$, obviously $A_{0}$ should be zero. Then the term with largest exponent of $t$ is $\frac{(-1)^{l_{1}}\mu_{0}}{\Gamma(1-l_{1}\alpha)}A_{l_{1}}$ $t^{-l_{1}\alpha}$. It follows that $A_{l_{1}}=0$. Still we can repeat the process to prove $A_{l_{k}}=0$ for all $k\in\mathbb{N}$. When $\mu_{0}=0$, we claim that $c_{\mu,j-1}$ are not identically zero. Then with the same method we can still prove $A_{l_{k}}=0$ for all $k\in\mathbb{N}$. We then apply Lemma 3.7 again to finish the proof.
It remains to prove that $\mu_{j}=0$ and $c_{\mu,m}=0$ for all $j,m\geq 0$ implies $\mu\equiv 0$. Since $\mu_{j}=0$ for all $j\geq 0$, by definition of $c_{\mu,m}$ we have $c_{\mu,m}=\int_{0}^{\infty}s^{m}\mu(s)\,\text{d}s$.
@@
So the condition (28) is satisfied. Since $c_{\mu,m}=0$ for all $m\geq 0$, the equation $\int_{0}^{\infty}p\mu^{(1)}\cdot\mu^{(2)}\,\text{d}s=0$ is true for all polynomials $p$. The decay property $|\mu^{(2)}(t)|\leq\exp(-c_{2}\sqrt{t}/2)$ implies that $\mu^{(2)}\in L^{1}(0,\infty)$, which further implies $\mu^{(2)}\,\text{d}s\in\left(C_{0}[0,\infty)\right)^{*}$. Finally we can use the density of the set $\{p\mu^{(1)}\,|\ p\text{ is a polynomial}\}$ to obtain $\mu^{(2)}\equiv 0$, which follows $\mu\equiv 0$. $\square$
@@
Now that $\alpha$ is irrational, if $-l_{k}\alpha-j=-l_{k^{\prime}}\alpha-j^{\prime}$, then $l_{k}=l_{k^{\prime}}$ and $j=j^{\prime}$. This shows that the terms in first two sums in equation (27) (or (29)) do not overlap. A comparison of coefficients yields $A_{l_{k}}\mu_{j}=\widetilde{A}_{l_{k}}\widetilde{\mu}_{j}$ for all $j\geq 0$ and $k\in\mathbb{N}$. It follows that $\sum_{n=1}^{\infty}(\mu_{j}a_{n}-\widetilde{\mu}_{j}\widetilde{a}_{n})\lambda_{n}^{-l_{k}-1}=0$. By Lemma 3.7, we have
@@
If $\mu_{j}=0$ for all $j\geq 0$, equation (30) implies that $\widetilde{\mu}_{j}=0$ for $j\geq 0$. Now, we can examine the expansions (27) and (29) again to deduce that $A_{l_{k}}c_{\mu,j}=\widetilde{A}_{l_{k}}c_{\widetilde{\mu},j}$ for $j\geq 0$ and $k\in\mathbb{N}$. Another application of Lemma 3.7 yields $c_{\mu,j}a_{n}-c_{\widetilde{\mu},j}\widetilde{a}_{n}=0$ for $j\geq 0$ and $n\in\mathbb{N}$. We have shown that there exists $j_{0}\geq 0$ such that $c_{\mu,j_{0}}\neq 0$ in the proof of Proposition 3.6. Now we can imitate the proof of the case $\mu_{j_{0}}\neq 0$ to obtain $a_{n}=\kappa\widetilde{a}_{n}$ for all $n\in\mathbb{N}$ and $\widetilde{\mu}=\kappa\mu$. $\square$

### v2

@@ 2 Preliminaries and main theorems
This section extends the discussion to equation (1). Define the domain of $A$ by $\mathcal{D}(A)=H^{2}(\Omega)\cap H_{0}^{1}(\Omega)$. Suppose $f\in L^{2}(\Omega)$ and $\mu\in L^{\infty}(0,\infty)$, then equation (1) admits a weak solution $u\in L^{2}_{\text{loc}}(0,\infty;$ $\mathcal{D}(A))\cap H^{\alpha}_{\text{loc}}(0,\infty;L^{2}(\Omega))$ as established in [19, Theorem 2.2]. Given that the spectrum of $A$ consists of positive eigenvalues of finite multiplicities, these can be denoted as $\{\lambda_{n}\}_{n\in\mathbb{N}}$ in increasing order, without counting multiplicities. Let $P_{n}$ be the orthogonal projection onto the eigenspace of $\lambda_{n}$. The solution $u$ is given by (refer to [22, Lemma 1])
@@
(a) $E_{\alpha,\beta}$ is entire in $\mathbb{C}$.
@@
$E_{\alpha,\beta}(-x)\sim\sum_{k=1}^{\infty}\frac{(-1)^{k+1}}{\Gamma(\beta-k\alpha)}x^{-k}\text{ as }x\rightarrow+\infty.$
Here $\frac{1}{\Gamma(z)}=0$ for $z=0,-1,-2,\ldots$.
@@
Then $||g||_{H^{2\sigma}(\Omega)}\leq C||A^{\sigma}g||_{L^{2}(\Omega)}$ for $g\in\mathcal{D}(A^{\sigma})$. Moreover, we denote $g(t)=O(t^{-\infty})$ if $g(t)=O(t^{-p})$ for any $p\in\mathbb{N}$.
@@
Assume that $\mu\in L^{\infty}(0,\infty)$ has an asymptotic expansion over $t^{-1}$:
@@
(a) Assume that $f\in L^{2}(\Omega)$ and
@@
(b) Let $f\in\mathcal{D}(A^{\sigma_{2}})$, where $\sigma_{2}>(d-1)/4$. Assume that
@@
Let $\mu\in L^{\infty}(0,\infty)$ satisfy the conditions (3) and (4a)/(4b) in Theorem 2.2. Define
@@
(a) Let $f\neq 0$ and $f\in\mathcal{D}(A^{\sigma_{3}})$, where $\sigma_{3}>\max\{0,d/2-1\}$. Then for any $x_{0}\in\Omega\setminus\Lambda(f)$,
@@
(b) Let $f\neq 0$ and $f\in\mathcal{D}(A^{\sigma_{4}})$, where $\sigma_{4}>d/2-1/4$. Then for any $x_{0}\in\partial\Omega\setminus\Lambda_{b}(f)$,
@@
Let $\alpha\in(0,2)\setminus\mathbb{Q}$. Suppose $u$ and $\widetilde{u}$ are solutions to the fractional diffusion-wave equation (1) with source terms $F(x,t)=f(x)\mu(t)$ and $\widetilde{F}(x,t)=\widetilde{f}(x)\widetilde{\mu}(t)$ respectively. Moreover, $\mu$ and $\widetilde{\mu}\in L^{\infty}(0,\infty)$ satisfy the conditions (3) and (4a)/(4b) in Theorem 2.2.
(a) Assume that $f,\widetilde{f}\in L^{2}(\Omega)$ and
@@
(b) Let $f,\widetilde{f}\in\mathcal{D}(A^{\sigma_{2}})$, where $\sigma_{2}>(d-1)/4$. Assume that
@@
Remark 2.3.
Our results cover the case when $\mu$ is compactly supported [5, 6, 22]. In this case $\mu$ has an asymptotic expansion of 00, and both assumptions (4a) and (4b) are satisfied. However, based on Laplace transform techniques in [5, 6], the authors proved the uniqueness with lower regularity conditions, namely, $f\in L^{2}(\Omega)$ for analogs of Theorems 2.2 and 2.4, and $f\in\mathcal{D}(A^{\sigma})$, $\sigma>(d-2)/4$ for an analog of Theorem 2.3(b).
@@ 3.1 The key propositions
$=C||\mu||_{L^{\infty}(0,t)}\,\lambda_{n}^{-1}\log(1+\lambda_{n})\left(1+\frac{\log((1+\lambda_{n})^{-1}+\lambda_{n}(1+\lambda_{n})^{-1}t^{\alpha})}{\log(1+\lambda_{n})}\right)$
@@
The last inequality follows from $\lambda_{n}\geq\lambda_{1}$. The proof is complete by treating the last term as a constant dependent on $t$. $\square$
@@ 3.2 The asymptotic expansion of $\psi_{n}(t)$
Since $\overline{R}_{\mu,m+1}(s)=O(s^{-m})$ as $s\rightarrow 0$ and $\overline{R}_{\mu,m+1}(s)=O(s^{-m-2})$ as $s\rightarrow\infty$, the quantity
$c_{\mu,m}:=\int_{0}^{\infty}s^{m}\overline{R}_{\mu,m+1}(s)\,\text{d}s$ (12)
is well defined.
@@
In the framework of Proposition 3.6, for any $N\in\mathbb{N}$, one can find $J$, $K$ and $M$ depending on $\alpha$ and $N$ only, such that $\psi_{n}(t)$ expands as follows:
@@
We emphasize that the bounding constant in the big-O notation is independent of $\lambda_{n}$ throughout the article.
Proof. Step 1: Treatment of the zeroth order term $\mu_{0}$.
Since $\psi_{n}$ is linear with respect to $\mu$, we first consider the zeroth order term $\mu_{0}$. Note that $E_{\alpha,1}(z)^{\prime}=\alpha^{-1}E_{\alpha,\alpha}(z)$ by definition, so
@@
Step 2: Expanding $\mu$ away from 0.
In the remaining proof, we set $\mu_{0}=0$ and introduce two parameters $\beta_{1}$ and $\beta_{2}$ so that $10<\beta_{1}<\beta_{2}<$ 1. Write $\mu(t)=\sum_{j=1}^{J}\mu_{j}t^{-j}+R_{\mu,J}(t)$, then $|R_{\mu,J}(t)|<Ct^{-J-1}$ for $t>1$. Since
@@
Step 3: Expanding $E_{\alpha,\alpha}(-\lambda_{n}s^{\alpha})$ for $s$ away from 0.
@@
with $R_{E,K}(\lambda_{n}s^{\alpha})=O\big{(}(\lambda_{n}s^{\alpha})^{-l_{K+1}-1}\big{)}$. We can choose $T_{1}$ sufficiently large and independent of $\lambda_{n}$ such that $t-t^{\beta_{1}}>\lambda_{n}^{-1/\alpha}t^{\beta_{2}}$ when $t>T_{1}$. This guarantees that, when $t>T_{1}$ and $0\leq s\leq t^{\beta_{1}}$, we have $\lambda_{n}(t-s)^{\alpha}>T_{1}^{\alpha\beta_{2}}$, which is independent of $n$. Together with $\mu\in L^{\infty}(0,\infty)$, we get
@@
Besides, we have $\lambda_{n}s^{\alpha}>T_{1}^{\alpha\beta_{2}}$ for $t>T_{1}$ and $\lambda_{n}^{-1/\alpha}t^{\beta_{2}}\leq s\leq t-t^{\beta_{1}}$. Hence, for each $j\in\mathbb{N}$,
@@
For $N\in\mathbb{N}$ given above, we choose $K=K(N)$ such that $\alpha\beta_{2}l_{K+1}>N$. Combining equations (14) – (17), we obtain
@@
Step 4: Expanding $(t-s)^{\gamma},\gamma<0$ in the integrals.
Since
$(1-s)^{\gamma}=\sum_{m=0}^{M}(-1)^{m}\binom{\gamma}{m}s^{m}+R_{\gamma,M}(s),$
we obtain that
@@
Step 5: Expanding integrals in the form of $\int s^{\gamma}F(s)\,\text{d}s$.
Step 5.1: Expansion of $\int_{0}^{t^{\beta_{1}}}s^{m}\mu(s)\,\text{d}s$.
We use the alternative expansion (10) of $\mu$ to expand the integral $\int_{0}^{t^{\beta_{1}}}s^{m}\mu(s)\,\text{d}s$. By equation (11), $\overline{R}_{\mu,J^{\prime}}=O(t^{-J^{\prime}-1})$ for $J^{\prime}>m$. Hence,
@@
$\int_{0}^{t^{\beta_{1}}}s^{m}\mu(s)\,\text{d}s$ $=\sum_{\begin{subarray}{c}1\leq j\leq J^{\prime}\\ j\neq m+1\end{subarray}}b_{j}^{\times}t^{(m-j+1)\beta_{1}}+\mu_{m+1}\log(1+t^{\beta_{1}})+c_{\mu,m}+O(t^{-N})$ (22)
$=\sum_{\begin{subarray}{c}1\leq j\leq J^{\prime}\\ j\neq m+1\end{subarray}}b_{j}^{\times}t^{(m-j+1)\beta_{1}}+\mu_{m+1}\beta_{1}\log t+c_{\mu,m}+O(t^{-N}).$
Step 5.2: Expansion of $\int_{0}^{t^{\alpha\beta_{2}}}s^{m/\alpha}E_{\alpha,\alpha}(-s)\,\text{d}s$.
We treat the integral $\int_{0}^{t^{\alpha\beta_{2}}}s^{m/\alpha}E_{\alpha,\alpha}(-s)\,\text{d}s$ similarly. When $m/\alpha\in\mathbb{N}\cup\{0\}$, observe that the coefficient of $s^{-m/\alpha-1}$ in the expansion of $E_{\alpha,\alpha}(-s)$ is 00, so for any $m\in\mathbb{N}\cup\{0\}$, we can use the ordinary expansion of $E_{\alpha,\alpha}$:
@@
$\int_{0}^{t^{\alpha\beta_{2}}}s^{m/\alpha}E_{\alpha,\alpha}(-s)\,\text{d}s=\sum_{k=1}^{K^{\prime}}b_{j,k}t^{(m-l_{k}\alpha)\beta_{2}}+b_{j,m}+O(t^{-N})$ (23)
for some $K^{\prime}=K^{\prime}(N)$ satisfying $(l_{K+1}\alpha-M)\beta_{2}>N$.
Step 5.3: Expansion of $\int_{\lambda_{n}^{-1/\alpha}t^{\beta_{2}-1}}^{1-t^{\beta_{1}-1}}s^{-l_{k}\alpha-1}(1-s)^{-j}\,\text{d}s$.
This integral can be expressed by the incomplete Beta function $B(a,b;x)$ and the hypergeometric function ${}_{2}\mathbf{F}_{1}(a,b,c;x)$ (see Section 8.17 and Chapter 15 of [16]). By equations (8.17.7) and (15.1.2) of [16] we obtain
@@
Step 6: Independence of the expansion from $\beta_{1}$ and $\beta_{2}$.
Finally we claim that the expansion of $\psi_{n}$ should be independent of $\beta_{1}$ and $\beta_{2}$ due to the uniqueness of asymptotic expansion. We choose $(\beta_{1},\beta_{2})$ and $(\beta_{1}^{\prime},\beta_{2}^{\prime})$ such that $\beta_{1}$, $\beta_{2}$, $\beta_{1}^{\prime}$, $\beta_{2}^{\prime}$ and $\alpha$ are algebraically independent over $\mathbb{Z}$. It follows that, if we pick a term in the first sum of the expansion of $\psi_{n}(t)$ with parameters $(\beta_{1},\beta_{2})$, the term still occurs under parameters $(\beta_{1}^{\prime},\beta_{2}^{\prime})$ if and only if the coefficients of $\beta_{1}$ and $\beta_{2}$ on the exponent of $t$ vanishes. Note that $l_{k}\alpha\notin\mathbb{N}\cup\{0\}$ by definition, so the only remaining term in the first sum is $\sum_{j=1}^{J^{\prime\prime}}\sum_{k=1}^{K^{\prime\prime}}b_{jk,j-1}^{\times}\lambda_{n}^{-l_{k}-1}t^{-l_{k}\alpha-j}$. This term can be brought into the second sum, again due to $l_{k}\alpha\notin\mathbb{N}\cup\{0\}$. $\square$
@@ 3.3 Proof of the key propositions
Lemma 3.10.
Suppose $\mu\in L^{\infty}(0,\infty)$ satisfies the conditions (3) and (4a)/(4b) in Theorem 2.2. Define $c_{\mu,m}$ according to equations (11) and (12). If $\mu_{j}=0$ and $c_{\mu,m}=0$ for all $j,m\geq 0$, then $\mu\equiv 0$.
Proof.
Since $\mu_{j}=0$ for all $j\geq 0$, by definition (11) and (12) of $c_{\mu,m}$ we have $c_{\mu,m}=\int_{0}^{\infty}s^{m}\mu(s)\,\text{d}s$.
@@
So the condition (28) is satisfied. Since $c_{\mu,m}=0$ for all $m\geq 0$, the equation $\int_{0}^{\infty}p\mu^{(1)}\cdot\mu^{(2)}\,\text{d}s=0$ is true for all polynomials $p$. The decay property $|\mu^{(2)}(t)|\leq\exp(-c_{2}\sqrt{t}/2)$ implies that $\mu^{(2)}\in L^{1}(0,\infty)$, which further implies $\mu^{(2)}\,\text{d}s\in\left(C_{0}[0,\infty)\right)^{*}$. Finally we can use the density of the set $\{p\mu^{(1)}\,|\ p\text{ is a polynomial}\}$ to obtain $\mu^{(2)}\equiv 0$, which follows $\mu\equiv 0$. ∎
$\sum_{n=1}^{\infty}|a_{n}\lambda_{n}^{-1}\log(1+\lambda_{n})|<\infty,$
By the decay condition (8), all the coefficients of $t^{\gamma}$ and $t^{\gamma}\log t$, $\gamma>-N$ in the expansion above (after collecting the terms) should be zero.
If there exists $j\geq 1$ such that $\mu_{j}\neq 0$, we choose the smallest such $j$ and denote it by $j_{0}$. We choose $N_{0}$ sufficiently large so that $J\geq j_{0}$ when $N\geq N_{0}$. In this situation we extract terms of the form $t^{\gamma}\log t$ from the expansion (29). Consequently, after collecting the terms, all the coefficients of $t^{-l_{k}\alpha-j}\log t$ in the expansion
$\sum_{j=j_{0}}^{J}\sum_{k=1}^{K}\frac{(-1)^{l_{k}-1+j}}{\Gamma(-l_{k}\alpha)}\binom{-l_{k}\alpha-1}{j-1}A_{l_{k}}\mu_{j}t^{-l_{k}\alpha-j}\log t,$ (30)
are zero as long as $l_{k}\alpha+j<N$. In this sum, the exponent $-l_{k}\alpha-j$ of $t$ takes the maximum if and only if $j=j_{0}$ and $k=1$. Since $l_{1}\alpha\notin\mathbb{N}\cup\{0\}$, we have $\frac{(-1)^{l_{1}-1+j_{0}}}{\Gamma(-l_{1}\alpha)}\neq 0$, which implies $A_{l_{1}}=0$. Now the summation over $k$ in expansion (30) starts from $k=2$, and the exponent $-l_{k}\alpha-j$ takes the maximum if and only if $j=j_{0}$ and $k=2$. Repeat the process above, and we can prove that $A_{l_{k}}=0$ for all $k$ such that $1\leq k\leq K$ and $l_{k}\alpha+j_{0}<N$. Letting $N\rightarrow\infty$, we obtain $A_{l_{k}}=0$ for all $k\in\mathbb{N}$. Finally the proposition is proved by applying Lemma 3.7 to $\{A_{l_{k}}\}_{k\in\mathbb{N}}$.
If $\mu_{0}\neq 0$, obviously $A_{0}$ should be zero. Then the term with largest exponent of $t$ is $\frac{(-1)^{l_{1}}\mu_{0}}{\Gamma(1-l_{1}\alpha)}A_{l_{1}}$ $t^{-l_{1}\alpha}$. It follows that $A_{l_{1}}=0$. Still we can repeat the process above to prove $A_{l_{k}}=0$ for all $k\in\mathbb{N}$. When $\mu_{0}=0$, then $c_{\mu,j-1}$ are not identically zero due to Lemma 3.10 and the assumption $\mu\not\equiv 0$. Using the same iteration process we can still prove $A_{l_{k}}=0$ for all $k\in\mathbb{N}$. We then apply Lemma 3.7 again to finish the proof. $\square$
@@
Now that $\alpha$ is irrational, so $-l_{k}\alpha-j=-l_{k^{\prime}}\alpha-j^{\prime}$ if and only if $l_{k}=l_{k^{\prime}}$ and $j=j^{\prime}$. This shows that the terms in first two sums in equation (29) (or (31)) do not overlap. A comparison of coefficients yields $A_{l_{k}}\mu_{j}=\widetilde{A}_{l_{k}}\widetilde{\mu}_{j}$ for all $j\geq 0$ and $k\in\mathbb{N}$. It follows that $\sum_{n=1}^{\infty}(\mu_{j}a_{n}-\widetilde{\mu}_{j}\widetilde{a}_{n})\lambda_{n}^{-l_{k}-1}=0$. By Lemma 3.7, we have
@@
If $\mu_{j}=0$ for all $j\geq 0$, equation (32) implies that $\widetilde{\mu}_{j}=0$ for $j\geq 0$. Now, we can examine the expansions (29) and (31) again to deduce that $A_{l_{k}}c_{\mu,j}=\widetilde{A}_{l_{k}}c_{\widetilde{\mu},j}$ for $j\geq 0$ and $k\in\mathbb{N}$. Another application of Lemma 3.7 yields $c_{\mu,j}a_{n}-c_{\widetilde{\mu},j}\widetilde{a}_{n}=0$ for $j\geq 0$ and $n\in\mathbb{N}$. According to the assumption $\mu\not\equiv 0$ and Lemma 3.10, there exists $j_{0}\geq 0$ such that $c_{\mu,j_{0}}\neq 0$. Now we can imitate the proof of the case $\mu_{j_{0}}\neq 0$ to obtain $a_{n}=\kappa\widetilde{a}_{n}$ for all $n\in\mathbb{N}$ and $\widetilde{\mu}=\kappa\mu$. $\square$